# Inclusions of real rank zero

### v2

@@ 5. Purely infinite inclusions
Let $a\in A_{+}$ be nonzero and $\epsilon>0$. It suffices to show that there exists a projection $p\in\overline{aBa}$ such that $p(a-\epsilon)_{+}=(a-\epsilon)_{+}$. We will construct this projection by obtaining a suitable scaling element in the sense of [5, Definition 1.1] (see also [37, Remark 2.4]). [^7] A contraction $x$ is called a scaling element if $x^{*}xxx^{*}=xx^{*}$. Following [37, Remark 2.4], to obtain such a projection $p$, it suffices to find a contraction $x\in\overline{aBa}$ such that
@@
Consider the spectrum of $a$ in $B$ denoted by $\sigma(a)$. If $\sigma(a)\cap(\epsilon/2,\epsilon)=\emptyset$, then we can take $p$ to be the orthogonal complement of the spectral projection of $a$ supported on $(\epsilon/2,\epsilon)$. Else, let $f:[0,\infty)\to[0,1]$ be a continuous function with open support $(\epsilon/2,\epsilon)$. Then, $f(a)$ is full and properly infinite in $B$, so $a\preceq f(a)$ ([32, Proposition 3.5]), with the Cuntz subequivalence taking place in the hereditary subalgebra $\overline{aBa}$.
Applying [37, Remark 2.5] to the subequivalence $a\preceq f(a)$, we can find a contraction $x\in\overline{aBa}$ such that
@@
It follows that $(x^{*}x)(xx^{*})=xx^{*}$, so $x$ is a scaling element in the sense of [5, Definition 1.1]. Moreover, as $(a-\epsilon)_{+}\in\overline{(a-\epsilon/2)_{+}B(a-\epsilon/2)_{+}}$, $x^{*}x(a-\epsilon)_{+}=(a-\epsilon)_{+}$. Furthermore, $xx^{*}(a-\epsilon)_{+}=0$, as $f(a)$ and $(a-\epsilon)_{+}$ are orthogonal. Following [37, Remark 2.4], $v=x+(1-x^{*}x)^{1/2}$ is an isometry in the unitisation of $\overline{aBa}$ such that $p=1-vv^{*}$ is a projection in $\overline{aBa}$ with $p(a-\epsilon)_{+}=(a-\epsilon)_{+}$. Hence $A\subseteq B$ has real rank zero.
@@
A particular, but important class of inclusions as in the statement of Theorem 5.1 is provided by the $*{}^{*}$-homomorphisms classified by the first named author in [23]. Therefore, we aim to obtain an equivalent characterisation for inclusions covered by Theorem 5.1.
Remark 5.3.
From the proof of Theorem 5.1, we see that whenever $\overline{aBa}$ is not unital i.e. $\{0\}\in\sigma(a)$ is not isolated, then the approximate unit of projections $(p_{n})_{n}$ in $\overline{aBa}$ may be chosen to be increasing, and such that each $p_{n}$ is full, properly infinite and with $[p_{n}]_{0}=0$ in $K_{0}(B)$.
In fact, the projection $p$ in the proof is of the form $p=1-vv^{\ast}$ for an isometry $v$ in the unitsiation of $\overline{aBa}$, and hence $[p]_{0}=0$ in $K_{0}(B)$.
@@
Theorem 5.1 shows that (i) implies (ii), so it remains to show that (ii) implies (i). Let $a\in A$ be a nonzero positive element and let $p\in\overline{aBa}$ be a full properly infinite projection. Then, $a$ is in the ideal generated by $p$, so $a\preceq p$ ([32, Proposition 3.5]). Since $p\in\overline{aBa}$, it also follows that $p\preceq a$ ([32, Proposition 2.7 (i)]), so $a$ is full. Moreover,
@@
Inclusions with infinite behaviour of this type have been essential in proving that various natural constructions are purely infinite. For example, in [45, Theorem 3.3], Rørdam and Sierakowski show that, under some assumptions, the crossed product $A\rtimes G$ is purely infinite if and only if any nonzero positive element in $A$ is properly infinite in $A\rtimes_{r}G$. Similarly, in [6], Bönicke and Li study when a groupoid $\mathrm{C}^{*}$-algebra is purely infinite. If $\mathcal{G}$ is a groupoid with unit space $\mathcal{G}^{(0)}$, they use the canonical inclusion $C_{0}(\mathcal{G}^{(0)})\subseteq\mathrm{C}^{*}_{r}(\mathcal{G})$ to describe when $\mathrm{C}^{*}_{r}(\mathcal{G})$ is purely infinite (see [6, Theorem B]).
Assuming nuclearity and $\mathcal{O}_{\infty}$-stability, we can strenghten the result in Theorem 5.1 and show that $*{}^{*}$-homomorphisms classified by the first named author in [23] are approximately factoring through a $\mathrm{C}^{*}$-algebra of real rank zero. In many cases of interest, this algebra can be assumed to be a Kirchberg algebra.
First, let us record some definitions. Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras with $A$ separable and let $\theta:A\to B$ be a $*{}^{*}$-homomorphism. Then, following [22, Definition 3.16], $\theta$ is said to be $\mathcal{O}_{\infty}$-stable if $\mathcal{O}_{\infty}$ embeds unitally in $B_{\infty}\cap\theta(A)^{\prime}/\operatorname{Ann}(\theta(A))$. [^8] Note that $\operatorname{Ann}(\theta(A))=\{b\in B_{\infty}:b\theta(A)+\theta(A)b=\{0\}\}$. Two $*{}^{*}$-homomorphisms $\phi,\psi:A\to B$, with $A$ separable, are approximately Murray-von Neumann equivalent if there exists $v\in B_{\infty}$ such that $v^{*}\phi(a)v=\psi(a)$ and $v\psi(a)v^{*}=\phi(a)$ for all $a\in A$.
@@
Let $A$ be a separable exact $\mathrm{C}^{*}$-algebra, $B$ be a $\sigma$-unital $\mathrm{C}^{*}$-algebra and $\theta:A\to B$ be a full, nuclear, $\mathcal{O}_{\infty}$-stable $*{}^{*}$-homomorphism. Then $\theta$ is approximately Murray-von Neumann equivalent to a $*{}^{*}$-homo-morphism which factors through a separable exact $\mathrm{C}^{*}$-algebra $C$ with real rank zero. Moreover, if $A$ is $\operatorname{KK}$-equivalent to a nuclear $\mathrm{C}^{*}$-algebra, [^9] For example if $A$ is nuclear or if it satisfies the $\operatorname{UCT}$. then $C$ can be taken to be a Kirchberg algebra. [^10] A simple, separable, nuclear, purely infinite $\mathrm{C}^{*}$-algebra is called a Kirchberg algebra.
@@
Building on the work of Pimsner in [38], Kumjian showed in [33] that there exists a simple separable purely infinite $\mathrm{C}^{*}$-algebra $C$ and an inclusion $\iota:A\to C$ such that $\iota$ is a $\operatorname{KK}$-equivalence. The existence of such $C$ is proved in [33, Theorem 3.1] (note that nuclearity is only needed to obtain nuclearity of $C$), while the fact that $\iota$ induces a $\operatorname{KK}$-equivalence follows from [38, Corollary 4.5]. Moreover, since $A$ is exact, $C$ can also be chosen to be exact by [19, Corollary 3.6].
Since the product of a $\operatorname{KK}$-element with a $\operatorname{KK}_{\operatorname{nuc}}$-element is in $\operatorname{KK}_{\operatorname{nuc}}$ ([47, Proposition 2.2(b)]), the product $\operatorname{KK}_{\operatorname{nuc}}(\theta)\circ\operatorname{KK}(\iota)^{-1}$ is an element of $\operatorname{KK}_{\operatorname{nuc}}(C,B)$. Then, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable $*{}^{*}$-homomorphism $\psi:C\to B$ such that $\operatorname{KK}_{\operatorname{nuc}}(\psi)=\operatorname{KK}_{\operatorname{nuc}}(\theta)\circ\operatorname{KK}(\iota)^{-1}$([23, Theorem $A$]). Now we apply the uniqueness part of the classification of full, nuclear, $\mathcal{O}_{\infty}$-stable $*{}^{*}$-homomorphisms. Precisely, since $\operatorname{KK}_{\operatorname{nuc}}(\psi\circ\iota)=\operatorname{KK}_{\operatorname{nuc}}(\theta)$, $\theta$ and $\psi\circ\iota$ are approximately Murray-von Neumann equivalent ([23, Theorem 8.10]). The fact that $C$ has real rank zero follows from [54] (alternatively apply Theorem 5.1 to the inclusion $C\subseteq C$).
Suppose now that $A$ is $\operatorname{KK}$-equivalent to a separable nuclear $\mathrm{C}^{*}$-algebra $D$. Let $x\in\operatorname{KK}(A,D)$ be an element inducing the equivalence. Note that since $D$ is nuclear, $\operatorname{KK}_{\operatorname{nuc}}(A,D)=\operatorname{KK}(A,D)$. Applying [33, Theorem 3.1] to the $\mathrm{C}^{*}$-algebra $D$, there exists a simple separable purely infinite $\mathrm{C}^{*}$-algebra $C$ and an inclusion $\iota:D\to C$ such that $\iota$ is a $\operatorname{KK}$-equivalence. Since $D$ is nuclear, it follows from [33, Theorem 3.1] that we can further assume that $C$ is nuclear, so $C$ is a Kirchberg algebra.
Then, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable $*{}^{*}$-homomorphism $\phi:A\to C$ such that
@@
([23, Theorem $A$]). Applying existence again, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable $*{}^{*}$-homomorphism $\psi:C\to B$ such that
@@
Since $\operatorname{KK}_{\operatorname{nuc}}(\psi\circ\phi)=\operatorname{KK}_{\operatorname{nuc}}(\theta)$, $\theta$ and $\psi\circ\phi$ are approximately Murray-von Neumann equivalent by [23, Theorem 8.10]. Hence, $\theta$ is approximately Murray-von Neumann equivalent to a $*{}^{*}$-homomorphism which factors through a Kirchberg algebra. ∎
@@
Let us denote this inclusion by $\iota$. If the group is non-amenable, there is an action $\alpha$ of $G$ on the Cantor space $X$ such that $C(X)\rtimes_{\alpha,r}G$ is a Kirchberg algebra ([45, Theorem 6.11]). Since the $\mathrm{C}^{*}$-algebra $C(X)\rtimes_{\alpha,r}G$ is simple and purely infinite, it has real rank zero by [54].
Suppose now that the group $G$ is amenable. Then, combining [48, Theorem 5.4] and [48, Lemma 5.2], there exists a totally disconnected compact Hausdorff space $X$ and a free minimal action $\alpha$ of $G$ on $X$ such that the transformation groupoid $X\rtimes_{\alpha}G$ is almost finite. [^11] All groupoids in [48] are assumed to be Hausdorff. Since the groupoid is also minimal and has totally disconnected unit space, the crossed product $C(X)\rtimes_{\alpha}G$ has real rank zero by [3, Theorem 3.2].  [^12] Note that in [3], minimal almost finite groupoids are assumed to be ample with compact unit space. In this case, ampleness is ensured by the fact that $X$ is totally disconnected and compact. Moreover, the definitions of almost finiteness in [3] and [48] coincide, as observed in the comments following [3, Remark 2.1].
We will now combine the two cases above. Let $X$ be a compact Hausdorff space and an action $\alpha$ of $G$ on $X$ such that the crossed product $C(X)\rtimes_{\alpha}G$ has real rank zero. Fix $x\in X$ and let $C(X)\to\ell^{\infty}(G)$ be the equivariant $*{}^{*}$-homomorphism given by $f\mapsto\left(f(\alpha_{g}(x))\right)_{g\in G}$. This induces a unital $*{}^{*}$-homomorphism $C(X)\rtimes_{\alpha,r}G\to\ell^{\infty}(G)\rtimes_{r}G$.
@@
Note that the inclusion $\mathrm{C}^{*}_{r}(G)\subseteq\ell^{\infty}(G)\rtimes_{r}G$ has real rank zero even if neither $\mathrm{C}^{*}_{r}(G)$ nor $\ell^{\infty}(G)\rtimes_{r}G$ have real rank zero. For example, take $G=\mathbb{F}_{2}\times\mathbb{Z}^{2}$ and note that its uniform Roe algebra $\ell^{\infty}(G)\rtimes G$ does not have real rank zero by [35, Corollary 3.5]. Moreover, $\mathrm{C}^{*}_{r}(G)\cong\mathrm{C}^{*}_{r}(\mathbb{F}_{2})\otimes C(\mathbb{T}^{2})$, which does not have real rank zero either, since its primitive ideal space is $\mathbb{T}^{2}$ (by simplicity of $\mathrm{C}^{*}_{r}(\mathbb{F}_{2})$ shown in [39]), which is not totally disconnected.
@@
If we denote the Gromov boundary of $G$ by $\partial G$, then there is an equivariant inclusion $C(\partial G)\subseteq C(\partial_{F}G)$ ([30, Remark 5.6]). Therefore, there is an induced unital $*{}^{*}$-homomorphism $C(\partial G)\rtimes_{r}G\to C(\partial_{F}G)\rtimes_{r}G$. It is readily seen that the composition
@@
We will now show that the crossed product $C(\partial G)\rtimes_{r}G$ is simple and purely infinite. Since the group $G$ is hyperbolic, the elements of $G$ are elliptic, parabolic, or hyperbolic ([25, Section 8.3]). Then, as $G$ is non-elementary and has no nontrivial finite normal subgroups, all elements of $G$ are hyperbolic ([25, Proposition 8.28, Theorem 8.29]). [^13] An element $g\in G$ is called hyperbolic if for any $x\in\partial G$, the orbit $n\in\mathbb{Z}\mapsto g^{n}x$ is a quasi-geodesic (see [25, Proposition 8.21]). Therefore, any element of $G$ fixes exactly two points of $\partial G$ ([25, Corollary 8.20]). Hence, the interior of fixed points is empty, so the action of $G$ on $\partial G$ is topologically free ([34, Definition 4]). Moreover, by [34, Example 2.1], the action of $G$ on $\partial G$ is a strong boundary action (in the sense of [34, Definition 1]). Hence, $C(\partial G)\rtimes_{r}G$ is simple and purely infinite by [34, Theorem 5]. In particular, $C(\partial G)\rtimes_{r}G$ has real rank zero by [54]. Thus, the conclusion follows. ∎
@@
We note what seems to be a persistent issue in the literature. In [1, Proposition 3.2], it is claimed that $C(\partial G)\rtimes_{r}G$ is simple and purely infinite if $G$ is a non-elementary hyperbolic group, while in [34], this is claimed to be the case for any word hyperbolic group. If $C(\partial G)\rtimes_{r}G$ is simple, then $G$ is a $\mathrm{C}^{*}$-simple group ([30, Theorem 1.5]). However, as it was pointed to us by Julian Kranz, there are non-elementary hyperbolic groups with finite normal subgroups (e.g. $\mathbb{F}_{2}\times\mathbb{Z}_{2})$ and in these cases, the group $G$ cannot be $\mathrm{C}^{*}$-simple (see for example [17, Theorem 8.14]). Thus, the assumption that $G$ has no nontrivial finite normal subgroups is necessary to conclude that the crossed product $C(\partial G)\rtimes_{r}G$ is simple.

### v3

@@ 5. Purely infinite inclusions
Let $a\in A_{+}$ be nonzero and $\epsilon>0$. It suffices to show that there exists a projection $p\in\overline{aBa}$ such that $p(a-\epsilon)_{+}=(a-\epsilon)_{+}$. We will construct this projection by obtaining a suitable scaling element in the sense of [5, Definition 1.1] (see also [39, Remark 2.4]). [^8] A contraction $x$ is called a scaling element if $x^{*}xxx^{*}=xx^{*}$. Following [39, Remark 2.4], to obtain such a projection $p$, it suffices to find a contraction $x\in\overline{aBa}$ such that
@@
Consider the spectrum of $a$ in $B$ denoted by $\sigma(a)$. If $\sigma(a)\cap(\epsilon/2,\epsilon)=\emptyset$, then we can take $p$ to be the orthogonal complement of the spectral projection of $a$ supported on $(\epsilon/2,\epsilon)$. Else, let $f:[0,\infty)\to[0,1]$ be a continuous function with open support $(\epsilon/2,\epsilon)$. Then, $f(a)$ is full and properly infinite in $B$, so $a\preceq f(a)$ ([34, Proposition 3.5]), with the Cuntz subequivalence taking place in the hereditary subalgebra $\overline{aBa}$.
Applying [39, Remark 2.5] to the subequivalence $a\preceq f(a)$, we can find a contraction $x\in\overline{aBa}$ such that
@@
It follows that $(x^{*}x)(xx^{*})=xx^{*}$, so $x$ is a scaling element in the sense of [5, Definition 1.1]. Moreover, as $(a-\epsilon)_{+}\in\overline{(a-\epsilon/2)_{+}B(a-\epsilon/2)_{+}}$, $x^{*}x(a-\epsilon)_{+}=(a-\epsilon)_{+}$. Furthermore, $xx^{*}(a-\epsilon)_{+}=0$, as $f(a)$ and $(a-\epsilon)_{+}$ are orthogonal. Following [39, Remark 2.4], $v=x+(1-x^{*}x)^{1/2}$ is an isometry in the unitisation of $\overline{aBa}$ such that $p=1-vv^{*}$ is a projection in $\overline{aBa}$ with $p(a-\epsilon)_{+}=(a-\epsilon)_{+}$. Hence $A\subseteq B$ has real rank zero.
@@
A particular, but important class of inclusions as in the statement of Theorem 5.1 is provided by the -homomorphisms classified by the first named author in [25]. Therefore, we aim to obtain an equivalent characterisation for inclusions covered by Theorem 5.1.
@@
From the proof of Theorem 5.1, we see that whenever $\overline{aBa}$ is not unital i.e. $0\in\sigma(a)$ is not isolated, then the approximate unit of projections $(p_{n})_{n}$ in $\overline{aBa}$ may be chosen to be increasing, and such that each $p_{n}$ is full, properly infinite and with $p_{n}]_{0}=0$ in $K_{0}(B)$.
In fact, the projection $p$ in the proof is of the form $p=1-vv^{\ast}$ for an isometry $v$ in the unitisation of $\overline{aBa}$, and hence $p]_{0}=0$ in $K_{0}(B)$.
@@
Theorem 5.1 shows that (i) implies (ii), so it remains to show that (ii) implies (i). Let $a\in A$ be a nonzero positive element and let $p\in\overline{aBa}$ be a full properly infinite projection. Then, $a$ is in the ideal generated by $p$, so $a\preceq p$ ([34, Proposition 3.5]). Since $p\in\overline{aBa}$, it also follows that $p\preceq a$ ([34, Proposition 2.7 (i)]), so $a$ is full. Moreover,
@@
Inclusions with infinite behaviour of this type have been essential in proving that various natural constructions are purely infinite. For example, in [47, Theorem 3.3], Rørdam and Sierakowski show that, under some assumptions, the crossed product $A\rtimes_{r}G$ is purely infinite if and only if any nonzero positive element in $A$ is properly infinite in $A\rtimes_{r}G$. Similarly, in [7], Bönicke and Li study when a groupoid $\mathrm{C}^{*}$-algebra is purely infinite. If $\mathcal{G}$ is a groupoid with unit space $\mathcal{G}^{(0)}$, they use the canonical inclusion $C_{0}(\mathcal{G}^{(0)})\subseteq\mathrm{C}^{*}_{r}(\mathcal{G})$ to describe when $\mathrm{C}^{*}_{r}(\mathcal{G})$ is purely infinite (see [7, Theorem B]).
Assuming nuclearity and $\mathcal{O}_{\infty}$-stability, we can strenghten the result in Theorem 5.1 and show that -homomorphisms classified by the first named author in [25] are approximately factoring through a $\mathrm{C}^{*}$-algebra of real rank zero. In many cases of interest, this algebra can be assumed to be a Kirchberg algebra.
First, let us record some definitions. Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras with $A$ separable and let $\theta:A\to B$ be a -homomorphism. Then, following [24, Definition 3.16], $\theta$ is said to be $\mathcal{O}_{\infty}$-stable if $\mathcal{O}_{\infty}$ embeds unitally in $B_{\infty}\cap\theta(A)^{\prime}/\operatorname{Ann}(\theta(A))$. [^9] Note that $\operatorname{Ann}(\theta(A))=\{b\in B_{\infty}:b\theta(A)+\theta(A)b=\{0\}\}$. Two -homomorphisms $\phi,\psi:A\to B$, with $A$ separable, are approximately Murray-von Neumann equivalent if there exists $v\in B_{\infty}$ such that $v^{*}\phi(a)v=\psi(a)$ and $v\psi(a)v^{*}=\phi(a)$ for all $a\in A$.
@@
Let $A$ be a separable exact $\mathrm{C}^{*}$-algebra, $B$ be a $\sigma$-unital $\mathrm{C}^{*}$-algebra and $\theta:A\to B$ be a full, nuclear, $\mathcal{O}_{\infty}$-stable -homomorphism. Then $\theta$ is approximately Murray-von Neumann equivalent to a -homo-morphism which factors through a separable exact $\mathrm{C}^{*}$-algebra $C$ with real rank zero. Moreover, if $A$ is $\operatorname{KK}$-equivalent to a nuclear $\mathrm{C}^{*}$-algebra, [^10] For example if $A$ is nuclear or if it satisfies the $\operatorname{UCT}$. then $C$ can be taken to be a Kirchberg algebra. [^11] A simple, separable, nuclear, purely infinite $\mathrm{C}^{*}$-algebra is called a Kirchberg algebra.
@@
Building on the work of Pimsner in [40], Kumjian showed in [35] that there exists a simple separable purely infinite $\mathrm{C}^{*}$-algebra $C$ and an inclusion $\iota:A\to C$ such that $\iota$ is a $\operatorname{KK}$-equivalence. The existence of such $C$ is proved in [35, Theorem 3.1] (note that nuclearity is only needed to obtain nuclearity of $C$), while the fact that $\iota$ induces a $\operatorname{KK}$-equivalence follows from [40, Corollary 4.5]. Moreover, since $A$ is exact, $C$ can also be chosen to be exact by [19, Corollary 3.6].
Since the product of a $\operatorname{KK}$-element with a $\operatorname{KK}_{\operatorname{nuc}}$-element is in $\operatorname{KK}_{\operatorname{nuc}}$ ([49, Proposition 2.2(b)]), the product $\operatorname{KK}_{\operatorname{nuc}}(\theta)\circ\operatorname{KK}(\iota)^{-1}$ is an element of $\operatorname{KK}_{\operatorname{nuc}}(C,B)$. Then, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable -homomorphism $\psi:C\to B$ such that $\operatorname{KK}_{\operatorname{nuc}}(\psi)=\operatorname{KK}_{\operatorname{nuc}}(\theta)\circ\operatorname{KK}(\iota)^{-1}$([25, Theorem $A$]). Now we apply the uniqueness part of the classification of full, nuclear, $\mathcal{O}_{\infty}$-stable -homomorphisms. Precisely, since $\operatorname{KK}_{\operatorname{nuc}}(\psi\circ\iota)=\operatorname{KK}_{\operatorname{nuc}}(\theta)$, $\theta$ and $\psi\circ\iota$ are approximately Murray-von Neumann equivalent ([25, Theorem 8.10]). The fact that $C$ has real rank zero follows from [56] (alternatively apply Theorem 5.1 to the inclusion $C\subseteq C$).
Suppose now that $A$ is $\operatorname{KK}$-equivalent to a separable nuclear $\mathrm{C}^{*}$-algebra $D$. Let $x\in\operatorname{KK}(A,D)$ be an element inducing the equivalence. Note that since $D$ is nuclear, $\operatorname{KK}_{\operatorname{nuc}}(A,D)=\operatorname{KK}(A,D)$. Applying [35, Theorem 3.1] to the $\mathrm{C}^{*}$-algebra $D$, there exists a simple separable purely infinite $\mathrm{C}^{*}$-algebra $C$ and an inclusion $\iota:D\to C$ such that $\iota$ is a $\operatorname{KK}$-equivalence. Since $D$ is nuclear, it follows from [35, Theorem 3.1] that we can further assume that $C$ is nuclear, so $C$ is a Kirchberg algebra.
Then, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable -homomorphism $\phi:A\to C$ such that
@@
([25, Theorem $A$]). Applying existence again, there exists a full, nuclear, $\mathcal{O}_{\infty}$-stable -homomorphism $\psi:C\to B$ such that
@@
Since $\operatorname{KK}_{\operatorname{nuc}}(\psi\circ\phi)=\operatorname{KK}_{\operatorname{nuc}}(\theta)$, $\theta$ and $\psi\circ\phi$ are approximately Murray-von Neumann equivalent by [25, Theorem 8.10]. Hence, $\theta$ is approximately Murray-von Neumann equivalent to a -homomorphism which factors through a Kirchberg algebra. ∎
@@
Let us denote this inclusion by $\iota$. If the group is non-amenable, there is an action $\alpha$ of $G$ on the Cantor space $X$ such that $C(X)\rtimes_{\alpha,r}G$ is a Kirchberg algebra ([47, Theorem 6.11]). Since the $\mathrm{C}^{*}$-algebra $C(X)\rtimes_{\alpha,r}G$ is simple and purely infinite, it has real rank zero by [56].
Suppose now that the group $G$ is amenable. Then, combining [50, Theorem 5.4] and [50, Lemma 5.2], there exists a totally disconnected compact Hausdorff space $X$ and a free minimal action $\alpha$ of $G$ on $X$ such that the transformation groupoid $X\rtimes_{\alpha}G$ is almost finite. [^12] All groupoids in [50] are assumed to be Hausdorff. Since the groupoid is also minimal and has totally disconnected unit space, the crossed product $C(X)\rtimes_{\alpha}G$ has real rank zero by [3, Theorem 3.2]. [^13] Note that in [3], minimal almost finite groupoids are assumed to be ample with compact unit space. In this case, ampleness is ensured by the fact that $X$ is totally disconnected and compact. Moreover, the definitions of almost finiteness in [3] and [50] coincide, as observed in the comments following [3, Remark 2.1].
We will now combine the two cases above. Let $X$ be a compact Hausdorff space and an action $\alpha$ of $G$ on $X$ such that the crossed product $C(X)\rtimes_{\alpha}G$ has real rank zero. Fix $x\in X$ and let $C(X)\to\ell^{\infty}(G)$ be the equivariant -homomorphism given by $f\mapsto\left(f(\alpha_{g}(x))\right)_{g\in G}$. This induces a unital -homomorphism $C(X)\rtimes_{\alpha,r}G\to\ell^{\infty}(G)\rtimes_{r}G$.
@@
Note that the inclusion $\mathrm{C}^{*}_{r}(G)\subseteq\ell^{\infty}(G)\rtimes_{r}G$ has real rank zero even if neither $\mathrm{C}^{*}_{r}(G)$ nor $\ell^{\infty}(G)\rtimes_{r}G$ have real rank zero. For example, take $G=\mathbb{F}_{2}\times\mathbb{Z}^{2}$ and note that its uniform Roe algebra $\ell^{\infty}(G)\rtimes G$ does not have real rank zero by [37, Corollary 3.5]. Moreover, $\mathrm{C}^{*}_{r}(G)\cong\mathrm{C}^{*}_{r}(\mathbb{F}_{2})\otimes C(\mathbb{T}^{2})$, which does not have real rank zero either, since its primitive ideal space is $\mathbb{T}^{2}$ (by simplicity of $\mathrm{C}^{*}_{r}(\mathbb{F}_{2})$ shown in [41]), which is not totally disconnected.
@@
If we denote the Gromov boundary of $G$ by $\partial G$, then there is an equivariant inclusion $C(\partial G)\subseteq C(\partial_{F}G)$ ([32, Remark 5.6]). Therefore, there is an induced unital -homomorphism $C(\partial G)\rtimes_{r}G\to C(\partial_{F}G)\rtimes_{r}G$. It is readily seen that the composition
@@
We will now show that the crossed product $C(\partial G)\rtimes_{r}G$ is simple and purely infinite. Since the group $G$ is hyperbolic, the elements of $G$ are elliptic, parabolic, or hyperbolic ([27, Section 8.3]). Then, as $G$ is non-elementary and has no nontrivial finite normal subgroups, all elements of $G$ are hyperbolic ([27, Proposition 8.28, Theorem 8.29]). [^14] An element $g\in G$ is called hyperbolic if for any $x\in\partial G$, the orbit $n\in\mathbb{Z}\mapsto g^{n}x$ is a quasi-geodesic (see [27, Proposition 8.21]). Therefore, any element of $G$ fixes exactly two points of $\partial G$ ([27, Corollary 8.20]). Hence, the interior of fixed points is empty, so the action of $G$ on $\partial G$ is topologically free ([36, Definition 4]). Moreover, by [36, Example 2.1], the action of $G$ on $\partial G$ is a strong boundary action (in the sense of [36, Definition 1]). Hence, $C(\partial G)\rtimes_{r}G$ is simple and purely infinite by [36, Theorem 5]. In particular, $C(\partial G)\rtimes_{r}G$ has real rank zero by [56]. Thus, the conclusion follows. ∎
@@
We note what seems to be a persistent issue in the literature. In [1, Proposition 3.2], it is claimed that $C(\partial G)\rtimes_{r}G$ is simple and purely infinite if $G$ is a non-elementary hyperbolic group, while in [36], this is claimed to be the case for any word hyperbolic group. If $C(\partial G)\rtimes_{r}G$ is simple, then $G$ is a $\mathrm{C}^{*}$-simple group ([32, Theorem 1.5]). However, as it was pointed out to us by Julian Kranz, there are non-elementary hyperbolic groups with finite normal subgroups (e.g. $\mathbb{F}_{2}\times\mathbb{Z}_{2})$ and in these cases, the group $G$ cannot be $\mathrm{C}^{*}$-simple (see for example [17, Theorem 8.14]). Thus, the assumption that $G$ has no nontrivial finite normal subgroups is necessary to conclude that the crossed product $C(\partial G)\rtimes_{r}G$ is simple.